OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –x – 2?

OpenStudy (anonymous):

y+x+2=0 having slope=-1 the line perpendicular eill have slope=-1/-1=1 thus using one point form: y-4=x-12